# Investigating Resource-efficient Neutron/Gamma Classification ML Models Targeting eFPGAs

### v1

@@ 3 Method
The neutron and gamma waveform data set used in this work was acquired with the same benchtop test bed described in detail in [16]. We briefly summarize the test bed again here. Figure 3, replicated from [16], provides a system diagram of the setup. A custom SiPM array fanout board was used to read out selected channels of an 8x8 array of OnSemi J-series 60035 SiPMs. Each SOUT channel is connected in series with a 50 resistor, which provides current-to-voltage conversion and appropriate impedance matching with the characteristic impedance of the coaxial connectors that the signal is routed to. Importantly, it also dominates the single photon response. The additional series resistance increases the decay constant associated with the recharging of the SiPM’s single photon avalanche diode (SPAD) microcells after receiving a hit. The waveform tails extend over a time period > 1 s as a result. This fact should be kept in mind when interpreting and extrapolating any of the results of this study. Each FOUT signal, already a voltage output, is connected to a balun transformer that provides the correct interface to the coaxial connector that brings the signal off the board. Signals are digitized using a commercial 14-bit, 250 MS/s CAEN DT5725 digitizer with data collected using a computer interface. Specifically, the digitized waveforms were collected for a 6mm cube of Stilbene optically coupled to a single SiPM in the array with AmBe used as a mixed neutron/gamma emitter.
In total, 80k gamma and 49k neutron waveforms were selected from a full 30 million event data set with an energy cut selecting only events with > 90 keVee energy depositions. The goal of this study was to demonstrate feasibility of ML model deployments on an eFPGA, not necessarily to improve classification efficiency at low energy depositions. The vast majority of events in the full data set correspond to the 60 keV Am-241 X-ray emission. Truth labels were assigned by performing a PSD ratio cut, with total integration set to 1500ns and partial integration set to 128ns, in the standard PSD ratio vs energy parameter space. Figure 4 shows these cuts and the resulting distribution of gamma and neutron events. The inversion of the neutron and gamma bands in the figure result from an inversion in the definition of PSD used. Instead of a tail-over-total, we performed a partial-over-total ratio as defined in Equation 3.1 with the start of the integration defined as , the partial integration ending at and the total at :
@@
Although the original waveforms were digitized with 14-bit resolution over 2V full-scale and centered at 0V at a sampling rate of 250 MS/s, as mentioned, we wanted to understand the effect of ADC bit-resolution and sampling rate on the performance of trained BDT and fcNN models. To do so, we can perform re-digitization, digital filtering and decimation of the waveform samples in software in order to emulate a range of bit-resolutions, front-end bandwidths and effective sampling rates. In total, we considered bit-resolutions from eight to the original 14 bits and sampling rates between 15.625 MS/s (decimation by a factor of 16) to the original 250 MS/s. In addition, re-filtering of the waveforms without downsampling was also performed to take a closer look at the expected degradation in performance of the benchmark charge comparison PSD method. It is important to note that the measured thermal noise of the test bed was 180V, which limited effective resolution to around 12 bits.
The CAEN digitizer uses a second-order, 125 MHz anti-aliasing filter. We emulated the filter in software using a second-order Butterworth filter implemented with the scipy Python package. The critical frequency (corner frequency) of the filter was varied to be half the effective targeted sampling rate to match the physical digitizer’s Nyquist configuration. Figure 5 shows a combined Bode plot of the various filter responses for the different targeted sampling rates. Five effective sampling rates were emulated by downsampling (every Nth sample kept) the original waveform samples by N, where N is the ratio of the original 250 MS/s sampling rate divided by the targeted sampling rate. In total, the parameter space encompassed by these two variables includes 35 unique pairs of effective bit-resolution and sampling rate. Figure 6 shows a representative FOUT and SOUT neutron waveform at the original 14 bit resolution and at a selection of other bit-resolutions down to 6 bits. No software re-filtering of the waveforms was performed for this figure.
In Section 2, a distinction was made on the degeneracy in physically implementing the extraction of input features. The two options discussed include using dedicated analog circuits and digitizing the final values or using a waveform digitizer and calculating input features digitally in a dedicated digital signal processing block. We can look at the performance of the standard charge comparison (PSD ratio) method under both scenarios. For both possibilities, 128ns was used for the partial integration window and 1024ns was used for the total integration window. The actual number of samples corresponding to these time windows varies depending on downsampling for the second option.
@@
We can plot the calculated partial and total integration values for both possible solutions. Figures 9, 10 plot the calculated partial and total integration values assuming an analog integration and digitization of final values, with the former figure showing an ideal 14-bit resolution on the final values and the latter assuming 10-bit resolution. The bottom panel for both figures show a zoomed out view of all events from the dataset used for the study and the top panel shows a zoomed in view of the lowest energy events above the 90 keVee equivalent cut. At 10-bit resolution, a significant degradation of the band separation is observed at the lowest energies compared to 14-bits, where band separation of the original waveforms is maintained across different effective bandwidths. It should be noted that a slight degradation at the lowest considered bandwidth is visible. On the other hand, Figures 11, 12 plot the results for the second option. We see that even at 14-bit resolution, band separation and widths worsen with decreasing effective bandwidth. At 8-bits resolution, a significant degradation of possible band separation is observed even at the highest bandwidth considered.
In general, it is clear that the standard charge comparison PSD metric is highly sensitive to the real-time implementation strategy and requires > 10 bit resolution and O(10) - O(100) MS/s for either possible implementation to maintain the ideal performance shown possible in [16]. In that paper, it was demonstrated that gamma leakages of < can be achieved while maintaining neutron efficiencies of 99.9%. Maintaining that performance equates to the ability to maintain as close to the ideal band separation and widths at the lowest energy deposition events considered in this work.
@@
For the feasibility study, we only consider the waveform digitizer based approach to input feature calculation. For ML models with multiple input features, a waveform digitizer and calculation of values for each feature digitally is, in general, less complex than designing analog circuits for each feature and offers a more flexible path towards re-definition, if required. BDTs and fcNNs have different model hyperparameters and, thus, considerations for resource efficiency. As a result, we independently constructed evaluation pipelines for these two ML architectures to account for such differences. These are discussed in the proceeding subsections. Regardless of the architecture, the same four input features were calculated for each waveform, for each of the 35 sampling rate, bit-resolution pairs. The four input features are the SOUT total integration value (integration window of 1024ns), SOUT partial integration value (integration window of 128ns), FOUT peak amplitude, and SOUT peak amplitude. The exact integration windows used were selected to maintain compatibility with the various corner frequencies used for re-filtering of waveforms. The total integration window was changed from that used for labeling truth data in an attempt to decouple any potential biases in model training. Specifically, the partial integration window was selected to ensure that the window encompassed a sufficient area past the waveform peak amplitude to perform PSD even at the lowest filter corner value. The total integration window was set large enough to integrate on the full tail of the waveform. The input features were settled on after a dedicated importance study of various options. The ease for either analog or digital calculation of the features was kept in mind.
The final study parameter is the base unit used for mapping ADC bit resolution to fixed point format. While it seems that using V as the base unit is the obvious choice to minimize input word length, the range of the input feature values and differences in ranges between the different input features plays a significant role in performance of fcNN models. For the BDT models, differences in model performance was also observed depending on choice of input feature value scale. For this reason, the choice was made to use mV as the base unit for both ML model architectures despite the larger number of input bits required. For the area input features used, an area unit of V*s was assumed so that range of values matched to an order of magnitude with the amplitude features. A more thorough investigation studying the effect of input feature value scale and range can be undertaken but is outside the scope of this paper. In addition, to provide optimal performance capability for the fcNN models, the input features were normalized to mean 0 and set to a standard deviation of 1. The scaler values to accomplish this can be extracted from the trained fcNN model to keep the same normalization for new input feature values. This will need to be implemented as a pre-processing step for an actual real-time execution of fcNN ML model inferences. Despite this, we kept the input word lengths the same across ML algorithms so that resource usage estimates can more closely correspond to differences from the two ML architectures.
@@ 3.1 Resource Efficient BDT Models
The terms and are unknown constants to be fitted for a given target, is the number of boosting rounds (num_rounds) and d is the max tree depth (max_depth). Interested readers can refer to [26] for plots that visually show the scaling resulting from this equation. Although the conifer implementation of the BDT allows selection of the quantization level for the input feature values, we assumed this was primarily determined by the ADC bit-resolution and appropriate scaling of the feature values as discussed. Similarly, the quantization level of the thresholds and the output raw scores were also selected respectively to account for the range of threshold values in the trained model consistent with the ranges of input feature values and to maintain overall performance of the model. Importantly, these parameters do not improve the performance of the BDT model past the baseline already set by the other parameters.
@@ 3.2 Resource Efficient fcNN Models
Much like XGBoost, QKeras for fcNNs has a large number of user configurable parameters. However, many more of the parameters have a non-trivial effect on the total resource usage of a trained model. In addition, unlike BDTs, there is no overall straightforward equation to describe the full parameter space’s relationship to total resource usage. This is partly because the logic synthesis tool can optimize the exact physical implementation of the circuit. This can obfuscate resource scaling relationships due to degeneracies in implementing certain operations with DSP or LUT resources. Optimally, the decision is left primarily to the logic synthesis software unless specified as a directive/pragma by the user. We can obtain expected scaling of resources if iterating on a single parameter while keeping all others constant, as should reasonably be expected. However, to obtain a resource-efficient final model, a defined and ordered procedure for selecting values for the many contributing hyperparameters need to be defined. In total, there are nine considered hyperparameters that can affect the total resource usage: input feature bit resolution, output probability bit resolution, bit resolution for weights, biases (in theory, this can be specified per hidden layer), bit resolution of the activation function, total number of (and size of each) hidden layers, and sparsity value target for pruning aware training. A quantized tanh activation function is assumed for the hidden layer(s) and a sigmoid activation function for the output layer.
@@
The order in which values are selected for these hyperparameters does matter. In general, the extent to which a model can be pruned/sparsified (i.e. the sparsity value target) without performance degredation partly depends on overall classification task complexity. Focusing on the neutron/gamma classification task at hand, it will depend strongly on the starting model size (number of hidden layers and size of each layer) as well as the input feature quantization level. So, we selected an ordered approach to defining values for the various hyperparameters.
@@
An initial look was conducted by using the unmodified waveforms recorded with the test bed (14 bits and 250 MS/s). The purpose was to narrow down the range of values for the overall hyperparameter space by finding lower bounds. From these initial results, a single hidden layer of size 8 was sufficient to ensure nearly unity neutron efficiency even with bit resolution for the weights and biases set to 6 bits and for the activation function set to 10 bits. The number of epochs was set to 120 after the standard optimization against rate of loss function minimization. A sparsity target value of up to 50% was found to still maintain neutron efficiency above 90% in the initial look. These starting values for the hyperparameters provide anchor points when introducing input feature bit resolution and shaping. Instead of a large number of possible values for the nine parameters, we only accept a small upwards perturbation from these anchor values. For example, number of hidden layers was only allowed to vary between 1 and 2 and the size of the hidden layers could be either 8 or 16 for each hidden layer. Similarly, the number of bits for the weights and biases was allowed to be either 6 or 8, and for the activation function, either 10 or 12 if needed. Sparsity values were allowed to vary between 10% to 70% and the number of epochs for training was fixed at 120. Unlike the number of boosting rounds parameter for BDTs, the number of epochs does not directly increase the model size.
@@ 4.1 BDT Results
We calculated the neutron efficiency value at a gamma leakage of 10-3 for each of 35 pairs of bit-resolution and re-filtered corner frequency values (with downsampling) at several (number of boosting rounds) values. At all bit-resolutions and corner frequencies, it is possible to achieve > 95% efficiency assuming an appropriate selection for . Even with = 50, we demonstrate that an ADC with only 8-bits resolution and Nyquist sampling rates of approximately 15 MS/s can still allow us to maintain excellent neutron efficiency at very low gamma leakage values in Figure 13. More generally, such a large parameter space for ADC specification where performance is maintained allows the final determination of design specs to be set by other system requirements for any specific target application.
@@
For testing purposes, a streamlined UART-based module to read/write to/from and execute ML inferences was implemented as a wrapper to the model implementation so that input words can be written to the FPGA dev board and results of an inference execution can be read back to verify performance on an actual Artix 7-series FPGA. The final nFF and nLUT values in the figure include the resources needed for implementing this module. A target clock frequency of 100MHz was used as a constraint. The required number of flip-flops is less than 3% of the total available on the target device. The required number of LUTs across the parameter space is less than 25% of the total available.
While an initial increase in nFFs with bit resolution is observed, the numbers level off at higher bit resolutions. This is not the case for nLUTs. The required nLUTs increase with bit resolution and for each bit resolution, with filter corner frequency. No digital signal processor (DSP) nor block RAM (BRAM) resources are required for a BDT model implementation. The breakdown of the hardware implementation of BDT models and the expected mapping to resource usage is given in [26]. The minute deviations from the expected trends are most likely a result of optimizations made during logic synthesis and physical implementation by the Vivado synthesis tool. Across all bit resolution and sampling rate pairs, initiation intervals of 1 clock cycle were achieved. Similarly, latencies between 3 to 4 clock cycles of were achieved. The clock period for the study was fixed at 10ns but in general can be lower, which would result in potentially lower absolute time values for initial intervals and latencies.
@@
We also calculated the neutron efficiency values at a gamma leakage of 10-3 for each of the 35 pairs of bit-resolution and re-filtered corner frequency values (with downsampling) for trained fcNN models. Because of the larger parameter space, no exact comparable parameter to for fcNNs is possible. A single hidden layer of size 16, with the number of bits used for the weights and biases equal to 8 and equal to 12 for the activation function, is sufficient to maintain high neutron efficiency values across all 35 considered pairs. This is true even at large sparsity target percentage values when performing pruning and quantization aware training. However, similar neutron efficiency values can also be obtained with a smaller hidden layer size and lower sparsity target values. We have used the larger size hidden layer model in Figure 15, which plots the observed statistical variance of fcNN models across a range of sparsity targets from 10% to 70%. As expected, with higher sparsity targets, the performance of the models at lower bit resolutions decreases, although overall neutron efficiency is still quite high. Significant statistical variance is not observed, unlike the BDT models. Some variance is seen at the largest sparsity values considered. Unlike the BDT models, the fcNN models were not able to maintain the same level of neutron efficiency performance at 8 bits across all filter corner values even at a minimal sparsity value target of 10%. Again, a total of 20 trials were used as a compromise between required computation time and having sufficient sample size to calculate values for each sparsity target value. Unlike for the BDT models, only one training iteration was performed per trial because the number of epochs is set sufficiently high with no significant increase in resource usages.
@@
Figure 16 histograms the number of flip-flops (nFFs) and number of look-up tables (nLUTs) for various bit resolution and corner frequency values as reported by Vivado post physical implementation at a sparsity target value of 30% for the above considered model. Like with the BDT model results, this value was chosen as representative of the edge of the plateau region. However, in this case, the decrease in neutron efficiency is minimal. The resources required to implement a fcNN model can be significantly reduced if the neutron efficiency target value is lowered, as is the case for the BDT models. We also see small fluctuations in the figure as a result of the synthesizer possibly optimizing out small parts of the model.
@@
The required number of flip-flops is less than 3% of the total available on the target device. The required number of LUTs across the parameter space is less than 10% of the total available. Unlike BDT models, BRAM and DSP resources are required to store the activation function in look-up table format and to perform the required multiplications when executing an inference. The number of DSPs are not explicitly plotted but they vary between two to 11 across the hyperparameter values defined for the figure. This is at maximum, 12% of the total available. Similarly, BRAM usage is reported as 0.5 of a single 18 Kb block for all models. This is the smallest unit reported and the true usage of the full block is expected to be different across the evaluated parameter space. An identical clock frequency constraint of 100MHz is specified. In addition, the same UART-based wrapper is used to read/write to/from and execute ML inferences to verify performance. Thus, resource usage numbers reported include this module as well and should be an identical constant number.
@@
Unlike the BDT models, resource usage for nFFs and nLUTs are relatively constant across the hyperparameter space specified. The exact reason can be extrapolated to how a hardware implementation is approached for the two ML architectures. Regardless, resources will scale with sparsity value target much like they do with for BDTs. The exact nFFs required for fcNN models are higher than for BDTs and the opposite is true for nLUTs. However, this needs to be considered against the additional DSPs and memory resource tiles required to implement fcNN models. The breakdown of the hardware implementation of fcNN models and general trends on resource usage is given in [27]. Across all bit resolution and sampling rate pairs, initiation intervals between 2 to 3 clock cycles were achieved. Similarly, latencies between 8 and 9 clock cycles were achieved. The BDT models clearly offer lower initiation intervals and latencies compared to fcNN models. This is expected from the constraints in gating the streaming of data between input to hidden and hidden to output layers for a given inference execution.

### v2

@@ 3 Method
The neutron and gamma waveform data set used in this work was acquired with the same benchtop test bed described in detail in [37]. We briefly summarize the test bed again here. Figure 3, replicated from [37], provides a system diagram of the setup. A custom SiPM array fanout board was used to read out selected channels of an 88 array of OnSemi J-series 60035 SiPMs [28]. Each standard, resistively coupled output (SOUT) signal is connected in series with a resistor, which provides current-to-voltage conversion and appropriate impedance matching with the characteristic impedance of the coaxial connectors that the signal is routed to. Importantly, it also dominates the single photon response. The additional series resistance increases the decay constant associated with the recharging of the SiPM’s single photon avalanche diode (SPAD) microcells after receiving a hit. The waveform tails extend over a time period > as a result. This fact should be kept in mind when interpreting and extrapolating any of the results of this study. Each fast, capacitively coupled output (FOUT) signal, already a voltage output, is connected to a balun transformer that provides the correct interface to the coaxial connector that brings the signal off the board. Signals are digitized using a commercial 14-bit, CAEN DT5725 digitizer with data collected using a computer interface. Specifically, the digitized waveforms were collected for a cube of Stilbene optically coupled to a single SiPM in the array with AmBe used as a mixed neutron/gamma emitter.
In total, 80k gamma and 49k neutron waveforms were selected from a full 30 million event data set with an energy cut selecting only events with > ee energy depositions. As stated, the goal of this study was to demonstrate feasibility of high performance ML model deployments on an eFPGA fabric, not necessarily to improve classification efficiency at low energy depositions. The vast majority of events in the full data set correspond to the Am-241 X-ray emission. Truth labels were assigned by performing a PSD ratio cut, with total integration set to and partial integration set to , in the standard PSD ratio vs energy parameter space. Figure 4 shows these cuts and the resulting distribution of gamma and neutron events. The inversion of the neutron and gamma bands in the figure result from an inversion in the definition of PSD used. Instead of a tail-over-total, we performed a partial-over-total ratio as defined in Equation (3.1) with the start of the integration defined as , the partial integration ending at and the total at :
@@
Although the original waveforms were digitized with 14-bit resolution over full-scale and centered at at a sampling rate of , as mentioned, we wanted to understand the effect of ADC bit-resolution and sampling rate on the performance of trained BDT and fcNN models. To do so, we can perform re-digitization, digital filtering and decimation of the waveform samples in software in order to emulate a range of bit-resolutions, front-end bandwidths and effective sampling rates. In total, we considered bit-resolutions from eight to the original 14 bits and sampling rates between (decimation by a factor of 16) to the original . In addition, re-filtering of the waveforms without downsampling was also performed to take a closer look at the expected degradation in performance of the benchmark charge comparison PSD method. It is important to note that the measured thermal noise of the test bed was , which limited effective resolution to around 12 bits.
The CAEN digitizer uses a second-order, anti-aliasing filter. We emulated the filter in software using a second-order Butterworth filter implemented with the SciPy [38] Python package. The critical frequency (corner frequency) of the filter was varied to be half the effective targeted sampling rate to match the physical digitizer’s Nyquist configuration. The corner frequency, also known as the point, represents the value at which the filter attenuates the input voltage signal by a factor of . Typically, the visual logarithmic representation of the frequency response is shown using a Bode plot [39]. Figure 5 shows this combined Bode plot of the various filter responses for the different targeted sampling rates. Five effective sampling rates were emulated by downsampling (every Nth sample kept) the original waveform samples by N, where N is the ratio of the original sampling rate divided by the targeted sampling rate. In total, the parameter space encompassed by these two variables includes 35 unique pairs of effective bit-resolution and sampling rate. Figure 6 shows a representative FOUT and SOUT neutron waveform at the original 14-bit resolution and at a selection of other bit-resolutions down to 6 bits. No software re-filtering of the waveforms was performed for this figure.
In Section 2, a distinction was made on the degeneracy in physically implementing the extraction of input features. The two options discussed include using dedicated analog circuits and digitizing the final values or using a waveform digitizer and calculating input features digitally in a dedicated digital signal processing block. We can look at the performance of the standard charge comparison (PSD ratio) method under both scenarios. For both possibilities, was used for the partial integration window and was used for the total integration window. The actual number of samples corresponding to these time windows varies depending on downsampling for the second option.
@@
We can plot the calculated partial and total integration values for both possible solutions. Figures 9, 10 plot the calculated partial and total integration values assuming an analog integration and digitization of final values, with the former figure showing an ideal 14-bit resolution on the final values and the latter assuming 10-bit resolution. The bottom panel for both figures show a zoomed out view of all events from the dataset used for the study and the top panel shows a zoomed in view of the lowest energy events above the ee equivalent cut. At 10-bit resolution, a significant degradation of the band separation is observed at the lowest energies compared to 14-bits, where band separation of the original waveforms is maintained across different effective bandwidths. It should be noted that a slight degradation at the lowest considered bandwidth is visible. On the other hand, Figures 11, 12 plot the results for the second option. We see that even at 14-bit resolution, band separation and widths worsen with decreasing effective bandwidth. At 8-bit resolution, a significant degradation of possible band separation is observed even at the highest bandwidth considered.
In general, it is clear that the standard charge comparison PSD metric is highly sensitive to the real-time implementation strategy and requires > 10-bit resolution and O(10) - O(100) MS/s for either possible implementation to maintain the ideal performance shown possible in [37]. In that paper, it was demonstrated that gamma leakages of < can be achieved while maintaining neutron efficiencies of 99.9%. Maintaining that performance equates to the ability to maintain as close to the ideal band separation and widths at the lowest energy deposition events considered in this work.
@@
For the feasibility study, we only consider the waveform digitizer based approach to input feature calculation. For ML models with multiple input features, a waveform digitizer and calculation of values for each feature digitally is, in general, less complex than designing analog circuits for each feature and offers a more flexible path towards re-definition, if required. BDTs and fcNNs have different model hyperparameters and, thus, considerations for resource efficiency. As a result, we independently constructed evaluation pipelines for these two ML architectures to account for such differences. These are discussed in the proceeding subsections. Regardless of the architecture, the same four input features were calculated for each waveform, for each of the 35 sampling rate, bit-resolution pairs. The four input features are the SOUT total integration value (integration window of ), SOUT partial integration value (integration window of ), FOUT peak amplitude, and SOUT peak amplitude. The exact integration windows used were selected to maintain compatibility with the various corner frequencies used for re-filtering of waveforms. The total integration window was changed from that used for labeling truth data in an attempt to decouple any potential biases in model training. Specifically, the partial integration window was selected to ensure that the window encompassed a sufficient area past the waveform peak amplitude to perform PSD even at the lowest filter corner value. The total integration window was set large enough to integrate on the full tail of the waveform. The input features were settled on after a dedicated importance study of various options. The ease for either analog or digital calculation of the features was kept in mind.
[…]
@@ 3.1 Resource Efficient BDT Models
The terms and are unknown constants to be fitted for a given target, is the number of boosting rounds (num_rounds) and is the max tree depth (max_depth). Interested readers can refer to [5] for plots that visually show the scaling resulting from this equation. Although the conifer implementation of the BDT allows selection of the quantization level for the input feature values, we assumed this was primarily determined by the ADC bit-resolution and appropriate scaling of the feature values as discussed. Similarly, the quantization level of the thresholds and the output raw scores were also selected respectively to account for the range of threshold values in the trained model consistent with the ranges of input feature values and to maintain overall performance of the model. Importantly, these parameters do not improve the performance of the BDT model past the baseline already set by the other parameters.
@@ 3.2 Resource Efficient fcNN Models
Much like XGBoost, QKeras for fcNNs has a large number of user configurable parameters. However, many more of the parameters have a non-trivial effect on the total resource usage of a trained model. In addition, unlike BDTs, there is no overall straightforward equation to describe the full parameter space’s relationship to total resource usage. This is partly because the logic synthesis tool can optimize the exact physical implementation of the circuit. This can obfuscate resource scaling relationships due to degeneracies in implementing certain operations with DSP or LUT resources. Optimally, the decision is left primarily to the logic synthesis software unless specified as a directive/pragma by the user. We can obtain expected scaling of resources if iterating on a single parameter while keeping all others constant, as should reasonably be expected. However, to obtain a resource-efficient final model, a defined and ordered procedure for selecting values for the many contributing hyperparameters need to be defined. In total, there are nine considered hyperparameters that can affect the total resource usage: input feature bit resolution, output probability bit resolution, bit resolution for weights, biases (in theory, this can be specified per hidden layer), bit resolution of the activation function, total number of (and size of each) hidden layers, and target sparsity values for pruning aware training. A quantized tanh activation function is assumed for the hidden layer(s) and a sigmoid activation function for the output layer. The importance of quantization and the need to incorporate sparsity and pruning methods are described in [36].
@@
The order in which values are selected for these hyperparameters does matter. In general, the extent to which a model can be pruned/sparsified (i.e. the target sparsity value) without performance degradation partly depends on overall classification task complexity. Focusing on the neutron/gamma classification task at hand, it will depend strongly on the starting model size (number of hidden layers and size of each layer) as well as the input feature quantization level. So, we selected an ordered approach to defining values for the various hyperparameters.
@@
An initial look was conducted by using the unmodified waveforms recorded with the test bed (14-bit and ). The purpose was to narrow down the range of values for the overall hyperparameter space by finding lower bounds. From these initial results, a single hidden layer of size 8 was sufficient to ensure nearly unity neutron efficiency even with bit-resolution for the weights and biases set to 6 bits and for the activation function set to 10 bits. The number of epochs was set to 120 after the standard optimization against rate of loss function minimization. A target sparsity value of up to 50% was found to still maintain neutron efficiency above 90% in the initial look. These starting values for the hyperparameters provide anchor points when introducing input feature bit resolution and shaping. Instead of a large number of possible values for the nine parameters, we only accept a small upwards perturbation from these anchor values. For example, number of hidden layers was only allowed to vary between 1 and 2 and the size of the hidden layers could be either 8 or 16 for each hidden layer. Similarly, the number of bits for the weights and biases was allowed to be either 6 or 8, and for the activation function, either 10 or 12 if needed. Target sparsity values were allowed to vary between 10% to 70% and the number of epochs for training was fixed at 120. Unlike the number of boosting rounds parameter for BDTs, the number of epochs does not directly increase the model size.
@@ 4.1 BDT Results
We calculated the neutron efficiency value at a gamma leakage of 10-3 for each of 35 pairs of bit-resolution and re-filtered corner frequency values (with downsampling) at several (number of boosting rounds) values. At all bit-resolutions and corner frequencies, it is possible to achieve > 95% efficiency assuming an appropriate selection for . Even with = 50, we demonstrate that an ADC with only 8-bit resolution and Nyquist sampling rates of approximately can still allow us to maintain excellent neutron efficiency at very low gamma leakage values in Figure 13. More generally, such a large parameter space for ADC specification where performance is maintained allows the final determination of design specs to be set by other system requirements for any specific target application.
@@
For testing purposes, a streamlined UART-based module was implemented as a wrapper to the model implementation so that input words can be written to the FPGA development board and results of an inference execution can be read back to verify performance on an actual Artix 7-series FPGA. The final nFF and nLUT values in the figure include the resources needed for implementing this module. A target clock frequency of was used as a constraint. The required number of flip-flops is less than 3% of the total available on the target device. The required number of LUTs across the parameter space is less than 25% of the total available.
While an initial increase in nFFs with bit resolution is observed, the numbers level off at higher bit resolutions. This is not the case for nLUTs. The required nLUTs increase with bit resolution and for each bit resolution, with filter corner frequency. No digital signal processor (DSP) nor block RAM (BRAM) resources are required for a BDT model implementation. The breakdown of the hardware implementation of BDT models and the expected mapping to resource usage is given in [5]. The minute deviations from the expected trends are most likely a result of optimizations made during logic synthesis and physical implementation by the Vivado synthesis tool. Across all bit resolution and sampling rate pairs, initiation intervals of 1 clock cycle were achieved. Similarly, latencies between 3 to 4 clock cycles of were achieved. The clock period for the study was fixed at but in general can be lower, which would result in potentially lower absolute time values for initial intervals and latencies.
@@
We also calculated the neutron efficiency values at a gamma leakage of 10-3 for each of the 35 pairs of bit-resolution and re-filtered corner frequency values (with downsampling) for trained fcNN models. Because of the larger parameter space, no exact comparable parameter to for fcNNs is possible. A single hidden layer of size 16, with the number of bits used for the weights and biases equal to 8 and equal to 12 for the activation function, is sufficient to maintain high neutron efficiency values across all 35 considered pairs. This is true even at large target sparsity values when performing pruning and quantization aware training. However, similar neutron efficiency values can also be obtained with a smaller hidden layer size and lower target sparsity values. We have used the larger size hidden layer model in Figure 15, which plots the observed statistical variance of fcNN models across a range of target sparsity values from 10% to 70%. As expected, with higher target sparsity values, the performance of the models at lower bit resolutions decreases, although overall neutron efficiency is still quite high. Significant statistical variance is not observed, unlike the BDT models. Some variance is seen at the largest target sparsity values considered. Unlike the BDT models, the fcNN models were not able to maintain the same level of neutron efficiency performance at 8 bits across all filter corner values even at a minimal target sparsity values of 10%. Again, a total of 20 trials were used as a compromise between required computation time and having sufficient sample size to calculate values for each target sparsity values. Unlike for the BDT models, only one training iteration was performed per trial because the number of epochs is set sufficiently high with no significant increase in resource usages.
@@
Figure 16 histograms the number of flip-flops (nFFs) and number of look-up tables (nLUTs) for various bit resolution and corner frequency values as reported by Vivado post physical implementation at a target sparsity values of 30% for the above considered model. Like with the BDT model results, this value was chosen as representative of the edge of the plateau region. However, in this case, the decrease in neutron efficiency is minimal. The resources required to implement a fcNN model can be significantly reduced if the neutron efficiency target value is lowered, as is the case for the BDT models. We also see small fluctuations in the figure as a result of the synthesizer possibly optimizing out small parts of the model.
@@
The required number of flip-flops is less than 3% of the total available on the target device. The required number of LUTs across the parameter space is less than 10% of the total available. Unlike BDT models, BRAM and DSP resources are required to store the activation function in look-up table format and to perform the required multiplications when executing an inference. The number of DSPs are not explicitly plotted but they vary between two to 11 across the hyperparameter values defined for the figure. This is at maximum, 12% of the total available. Similarly, BRAM usage is reported as 0.5 of a single block for all models. This is the smallest unit reported and the true usage of the full block is expected to be different across the evaluated parameter space. An identical clock frequency constraint of is specified. In addition, the same UART-based wrapper is used to send input words to the FPGA development board and read back the executed inference result to verify performance. Thus, resource usage numbers reported include this module as well and should be an identical constant number.
@@
Unlike the BDT models, resource usage for nFFs and nLUTs are relatively constant across the hyperparameter space specified. The exact reason can be extrapolated to how a hardware implementation is approached for the two ML architectures. Regardless, resources will scale with target sparsity values much like they do with for BDTs. The exact nFFs required for fcNN models are higher than for BDTs and the opposite is true for nLUTs. However, this needs to be considered against the additional DSPs and memory resource tiles required to implement fcNN models. The breakdown of the hardware implementation of fcNN models and general trends on resource usage is given in [2]. Across all bit resolution and sampling rate pairs, initiation intervals between 2 to 3 clock cycles were achieved. Similarly, latencies between 8 and 9 clock cycles were achieved. The BDT models clearly offer lower initiation intervals and latencies compared to fcNN models. This is expected from the constraints in gating the streaming of data between input to hidden and hidden to output layers for a given inference execution.